OpenStudy (anonymous):

Omar throws a softball from a starting height of 8 feet, with an initial velocity of 18 feet per second. Which of the following equations shows the path of the softball? H(t) = −16t2 + 26 H(t) = −16t2 + 26t H(t) = −16t2 + 8t + 18 H(t) = −16t2 + 18t + 8

OpenStudy (welshfella):

note at time = 0 h = 8

OpenStudy (anonymous):

The starting height is the height at time zero. So you need to choose a function with H(0) = 8. And the velocity is the derivative of position, so you need the function with H'(0) = 18. What do you think?

OpenStudy (anonymous):

i think the answer is A

OpenStudy (anonymous):

If you put t=0 into A, do you get H(0) = 8 ?

OpenStudy (welshfella):

you get 26 if you put t = 0 into A

OpenStudy (anonymous):

just kidding its C

OpenStudy (anonymous):

Right on. Well done!

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

put t = 0 into C and what do you get?

OpenStudy (anonymous):

You're welcome.

OpenStudy (welshfella):

?????

OpenStudy (welshfella):

Its D

OpenStudy (welshfella):

h is 8 when t = 0 h(0) = -16(0)^2 + 18(0) = 8 = 8

OpenStudy (anonymous):

plug O in for T and you get C

OpenStudy (anonymous):

were not trying to get 8 were trying to get 18

OpenStudy (welshfella):

plug 0 for t in C and you get h = 18 - but its not 18 at time 0 its 8

OpenStudy (welshfella):

believe me the answer is D

OpenStudy (welshfella):

Omar is not 18 feet tall!! lol

OpenStudy (anonymous):

were trying to find the path of the softball

OpenStudy (welshfella):

I know that but if you want to accept the wrong answer its up to you